The population mean weight of a candy is 10 grams with a standard deviation of 2 grams. A sample of 100 candies is taken. What is the probability that the sample mean is 9.7 grams or lower?
1
Expert's answer
2020-10-25T18:38:35-0400

Given : "\mu=10\\ \sigma=2"

n=100


Formula: "Z= \frac{\overline x-\mu}{\frac{\sigma}{\sqrt{n}}}"

Solution:

substitute the values to find z

"Z= \frac{9.7-10}{\frac{2}{\sqrt{100}}}" −1.5


Using z table,"Pr( \overline x≤9.7)=Pr(Z≤−1.5)=0.0668"